# An MLIR-based Compilation Framework for Control Flow Management on CGRAs

### v1

@@ 9.4. Comparison with state-of-the-art methodologies
Our compilation approach achieves the lowest compilation runtime across all benchmarks, as shown in Figure 14. We attain average speedups of 2.12×, 1.48×, and 1.20× over EffiMap (pc_control_mapping, ), SAT-MapIt (satjournal, ), and Marionette (control_plane, ), respectively. EffiMap does not fully exploit the optimization spaces during the compilation process: it relies on dedicated hardware to support instruction-level parallelism, but neglects loop-level parallelism opportunities. However, loop-level parallelism is critical for optimizing runtime efficiency for CGRA, as discussed in Section 9.3, where modulo scheduling achieves $L\approx 2*II$. Our CFG adaptation creates new kernels with loop length $L^{\prime}=II\approx 0.5*L$ on average. EffiMap achieves faster runtime for FIR because modulo scheduling cannot be performed due to the presence of non-mergeable branches inside an innermost loop. Our compilation method outperforms theirs because their heuristic strategy does not optimize for the shortest basic block runtime, leading to suboptimal solutions.
@@
When the computational workloads dominate the benchmarks with control-free, non-nested loops, both SATMapIt and Marionette achieve performance comparable to our method, as the runtime reconfiguration overhead becomes negligible. This is shown in benchmarks such as gsm and sha. However, both SatMapIt and Marionette experience significant reconfiguration overhead for nested loops. For example, a conv3d application with an input tensor of dimensions $D\times H\times W$ (depth × height × width) and a kernel of dimensions $d\times h\times w$ would require reconfiguration $D\times H\times W\times d\times h$ times–the number of branches to execute the innermost loop over the kernel width $w$. Even if the reconfiguration time is optimized to take only a few clock cycles, the large number of reconfigurations still introduces a substantial overhead. For conv2d and conv3d, the computation time is faster for Marionette and SatMapIt than for our approach, as in these methods, no live-in / live-out values must be maintained across basic blocks. Nonetheless, our approach still outperforms both when reconfiguration time is accounted for, because the entire benchmark can be mapped as a single computational kernel.
In conclusion, our compilation framework outperforms methodologies that rely on hardware reconfiguration units by eliminating the need for reconfiguration during basic block transitions. It also surpasses other PC-based control models through CFG reshaping, which enables effective support for modulo scheduling.
@@ 10. Conclusion
By combining efficiency and reprogrammability, CGRAs can support the acceleration of computationally intensive edge applications. The key to their mainstream adoption is the availability of a compilation framework enabling the automatization of the deployment of complex compute kernels on available hardware resources.
Against this backdrop, we have proposed in this paper a novel end-to-end framework that, by leveraging and extending the MLIR infrastructure, is able to generate CGRA assembly from source code. We broadened the scope of CGRA scheduling by considering control flow as well as data flow in the compilation process, showcasing how this approach can be employed to efficiently compile kernels. Moreover, our framework supports compilation for applications with arbitrary control flows without dedicated hardware support. The introduced CFG-based optimizations lead to speedups of 1.96x and 2.23x on 3x3 and 4x4 CGRAs, respectively, resulting in, on average, 2.12x faster executions compared to state-of-the-art methods.

### v2

@@ 9.4. Comparison with state-of-the-art methodologies
Our compilation approach achieves the lowest compilation runtime across all benchmarks, as shown in Figure 14. We achieve average speedups of 2.12×, 1.48×, and 1.20× over EffiMap (pc_control_mapping, ), SAT-MapIt (satjournal, ), and Marionette (control_plane, ), respectively. EffiMap does not fully exploit the optimization spaces during the compilation process: it relies on dedicated hardware to support instruction-level parallelism, but neglects loop-level parallelism opportunities. However, loop-level parallelism is critical for optimizing runtime efficiency for CGRA, as discussed in Section 9.3, where modulo scheduling achieves $L\approx 2*II$. Our CFG adaptation creates new kernels with loop length $L^{\prime}=II\approx 0.5*L$ on average. EffiMap achieves faster runtime for FIR because modulo scheduling cannot be performed due to the presence of non-mergeable branches inside an innermost loop. Our compilation method outperforms theirs because their heuristic strategy does not optimize for the shortest basic block runtime, leading to suboptimal solutions.
@@
When the computational workloads dominate the benchmarks with control-free, non-nested loops, both SATMapIt and Marionette achieve performance comparable to our method, as the runtime reconfiguration overhead becomes negligible. This is shown in benchmarks such as gsm and sha. However, both SatMapIt and Marionette experience significant reconfiguration overhead for nested loops. For example, a conv3d application with an input tensor of dimensions $D\times H\times W$ (depth × height × width) and a kernel of dimensions $d\times h\times w$ would require reconfiguration $D\times H\times W\times d\times h$ times, the number of branches to execute the innermost loop over the width of the kernel $w$. Even if the reconfiguration time is optimized to take only a few clock cycles, the large number of reconfigurations still introduces a substantial overhead. For conv2d and conv3d, the computation time is faster for Marionette and SatMapIt than for our approach, as in these methods, no live-in / live-out values must be maintained across basic blocks. Nevertheless, our approach still outperforms both when reconfiguration time is accounted for, because the entire benchmark can be mapped as a single computational kernel.
In conclusion, our compilation framework outperforms methodologies that rely on hardware reconfiguration units by eliminating the need for reconfiguration during basic block transitions. It also surpasses other PC-based control models through CFG reshaping, enabling effective support for modulo scheduling.
@@ 10. Conclusion
By combining efficiency and reprogrammability, CGRAs can support acceleration of computationally intensive edge applications. The key to their mainstream adoption is the availability of a compilation framework that enables the automatization of the deployment of complex compute kernels on available hardware resources.
Against this backdrop, we have proposed in this paper a novel end-to-end framework that, by leveraging and extending the MLIR infrastructure, is able to generate CGRA assembly from source code. We broadened the scope of CGRA scheduling by considering both control flow and data flow in the compilation process, showcasing how this approach can be employed to efficiently compile kernels. Moreover, our framework supports the compilation for applications with arbitrary control flows without dedicated hardware support. The introduced CFG-based optimizations lead to speedups of 1.96x and 2.23x on 3x3 and 4x4 CGRAs, respectively, resulting in, on average, 2.12x faster executions compared to state-of-the-art methods.
Acknowledgment
This work has been partially supported by the Swiss NSF Edge-Companions project (GA No. 10002812) and by the Swiss State Secretariat for Education, Research, and Innovation (SERI) through the SwissChips research project. This research was partially conducted by ACCESS – AI Chip Center for Emerging Smart Systems, supported by the InnoHK initiative of the Innovation and Technology Commission of the Hong Kong Special Administrative Region Government. This work was supported by the Swiss National Science Foundation via project ADApprox (grant 200020_188613).